### v1

# Causal Inference with Missing Exposures, Missing Outcomes, and Dependence

Abstract
Missing data are ubiquitous in public health research. The missing-completely-at-random (MCAR) assumption is often unrealistic and can lead to meaningful bias when violated. The missing-at-random (MAR) assumption tends to be more reasonable, but guidance on conducting causal analyses under MAR is limited when there is missingness on multiple variables. We present a series of causal graphs and identification results to demonstrate the handling of missing exposures and outcomes in observational studies. For estimation and inference, we highlight the use of targeted minimum loss-based estimation (TMLE) with Super Learner to flexibly and robustly address confounding, missing data, and dependence. Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed notable challenges due to confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), missingness on the outcome at follow-up (capturing who acquired TB), and clustering within households. Application to real data from SEARCH-TB highlighted the real-world consequences of the discussed methods. Estimates from TMLE suggested that alcohol use was associated with a 49% increase in the relative risk (RR) of incident TB infection (RR=1.49, 95%CI: 1.39–1.59). These estimates were notably larger and more precise than estimates from inverse probability weighting (RR=1.13, 95%CI: 1.00–1.27) and unadjusted, complete case analyses (RR=1.18, 95%CI: 0.89–1.57). Our work demonstrates the utility of causal models for describing the missing data mechanism and TMLE for flexible inference.
Keywords: Causal Inference, Missing Data, Targeted Minimum Loss-based Estimation (TMLE), Ensemble Machine Learning, Super Learner, Tuberculosis, Missing at Random
1 Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. A common approach is to simply exclude observations with missing data on the relevant variables. This “complete case” analysis can provide unbiased estimates if the data are missing as a result of a completely random process (i.e., missing-completely-at-random [MCAR]) [9]. However, this assumption is often unrealistic in health settings, leading to biased estimates as well as reduced statistical power (e.g., [10]). The Missing-at-Random (MAR) assumption allows missingness to depend on observed study variables and is often more plausible [9].
Suppose, for example, we aim to estimate the prevalence of a disease in the target population. On $N$ randomly sampled participants, we measure baseline covariates (e.g., age, gender, socio-economic status), but do not fully ascertain the outcome of interest. To recover the population-level outcome prevalence under MAR, we would need that the covariates capture all the common causes of measurement and outcomes as well as a positive probability of measurement within all possible covariate values. Then using standardization (a.k.a., G-computation), inverse weighting, multiple imputation, or more robust methods, we could obtain a point estimate and 95% confidence intervals using data on all $N$ participants. Instead, if we were willing to consider the stronger MCAR assumption, corresponding to no common causes of measurement and the outcome, then we could estimate disease prevalence with an empirical proportion among those measured.
This prevalence example illustrates the consequences of causal assumptions on missing data: stronger assumptions allow for simpler estimation approaches, while weaker assumptions require more complex estimation approaches. An analogous scenario arises in analyses of observational data, which are subject to confounding. In the following, we present a comprehensive framework to address common real-world challenges: confounding, missingness on multiple variables, and dependence between participants in our study. Our work builds on a long history of methods to address missing data (e.g., [11, 9, 12, 13, 14, 15]). As detailed below, we give special consideration to missing exposures. To the best of our knowledge, methods to handle missing exposure have largely focused on case-control designs, where exposure information is retrospectively collected on a sample of participants based on their outcome status (e.g., [16, 17, 4, 18, 19, 20]). Here, we focus on standard cross-sectional or longitudinal studies with missing exposures.
We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our target population would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential measurement of outcomes at baseline. Using Counterfactual Strata Effects [21, 5, 22, 23, 24, 25, 26], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios. Finally, we provide extensions for settings where participants are not independent and, instead, clustered within households or communities. Altogether, our presentation covers multi-source missing data, confounding, and dependence. We build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers. We illustrate the practical relevance with the SEARCH-TB study.
2 Motivating example
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [27]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [28]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection, a proxy for population-level transmission, was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [29, 30]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we provide an in-depth exploration of the methods to evaluate the effect of alcohol use on incident TB infection in SEARCH-TB [31]: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population of interest, (3) missingness in the final outcome, and (4) confounding. Additionally, given our focus on infectious disease outcomes, the independent and identically distributed (i.i.d.) assumption was likely violated, and we accounted dependence among study participants.
3 Related Causal Problems: Building Complexity
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or the causal risk ratio).
3.1 Classic point-treatment problem
First, we consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acylic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 1 and reflect the simplifying assumption that participants are independent (e.g., cannot transmit TB to one another due to geographical distance).
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
Figure 1: Directed acyclic graph (DAG) for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. The corresponding non-parametric structural equation model (NPSEM) is given by $L=f_{L}(U_{L})$; $A=f_{A}(L,U_{A})$; $Y=f_{Y}(L,A,U_{Y})$ where $(U_{L},U_{A},U_{Y})$ represent unmeasured factors determining the values of the covariates, exposure, and outcome, respectively. On the DAG, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$ and can be represented as $Y^{a}\perp A\mid L$. Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e.. Under these two assumptions, our causal target is equal to the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [12]. Even if these assumptions are not reasonable (e.g., there are unmeasured confounders in Figure 1), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
3.2 Missing exposures
Most real-world studies, however, depart from this idealized point-treatment problem. The first departure we consider is missingness on the exposure of interest. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is not observed (i.e., equal to NA). The DAG and NPSEM for such a study are given in Figure 2. To define causal effects when the exposure is subject to missingness, we now consider counterfactuals indexed by both the exposure and its measurement indicator. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and was at level $A=a$.
Figure 2: DAG for a point-treatment problem with missingness on the exposure: $L$=baseline covariates, $\Delta_{A}$=indicator of exposure measurment, $A$=exposure, $Y$=outcome, and $U^{*}$=unmeasured common causes of at least two variables in $(L,\Delta_{A},A,Y)$. The corresponding NPSEM is $L=f_{L}(U_{L})$; $\Delta_{A}=f_{\Delta_{A}}(L,U_{\Delta_{A}})$, $A=f_{A}(L,\Delta_{A},U_{A})$, $Y=f_{Y}(L,\Delta_{A},A,U_{Y})$.
Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had their exposure measured $\Delta_{A}=1$ and it were at level $A=a$. To identify this causal parameter and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we would need $L$ to be sufficient to control for confounding and differential measurement. In more detail, the following assumptions are required:
  • Within values of the covariates $L$, outcomes among those with measured exposures are representative of outcomes among those without measured exposures: $Y^{*}\perp\Delta_{A}\mid L$.
  • There is a non-zero probability of exposure measurement within all possible values of $L$: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e..
  • Among those with measured exposures ($\Delta_{A}=1$),
    • $L$ captures all the common causes of the exposure and outcome: $Y^{*}\perp A\mid\Delta_{A}=1,L$.
    • there is a non-zero probability of being exposed to level $A=a$ within all possible values of $L$: $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in eAppendix A. As before, even if these assumptions do not hold (e.g., there are unmeasured common causes of measurement and the outcome in Figure 2), we still have a well-defined statistical estimand for estimation and inference.
3.3 Missing Exposures and Outcomes
We now add another common complication: missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates, $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment, and $\Delta_{Y}$ be an indicator of outcome measurement. Specifically, if $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as before. However, if $\Delta_{Y}=0$ for a participant, their outcome $Y$ is not observed (i.e., equal to NA). The simplified DAG and full NPSEM for such a study are given in Figure 3.
Figure 3: DAG with missingness on the exposure and outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, $A$=exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of measured outcome, and $Y$=outcome. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least two variables in $(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,U_{L_{1}})$; $\Delta_{Y}=f_{\Delta_{Y}}(L_{0},\Delta_{A},A,L_{1},U_{\Delta_{Y}})$; $Y=f_{Y}(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},U_{Y})$.
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization and find a set of covariates that satisfies the backdoor criteria for each “intervention” node given the observed past [12]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$),
  • the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in eAppendix B [12, 32, 33]. As before, even if these identification assumptions do not hold, we still have a well-defined statistical estimand to focus our estimation efforts.
3.4 Missing Exposures and Outcomes at Baseline and Follow-up
We now consider missingness on the outcome at baseline. Continuing our running example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our cohort of participants without TB at baseline is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at baseline, $Y_{0}$ be an indicator having the outcome at baseline, $\Delta_{Y_{1}}$ be an indicator of outcome measurement at follow-up, and $Y_{1}$ be an indicator having the outcome at follow-up. As previously introduced, the value of the baseline outcome defines the target population; specifically, we are interested in occurrence of the outcome ($Y_{1}=1$) among those who are at risk at baseline ($Y_{0}=0$). The corresponding causal models can be found in Figure 4.
Figure 4: DAG with missingness on the exposure, the baseline outcome, and the follow-up outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, A=exposure, $\Delta_{Y_{0}}$=indicator of measured baseline outcome, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of measured outcome at follow-up, and $Y_{1}$= outcome at follow-up. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least 2 variables in $(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $\Delta_{Y_{0}}=f_{\Delta_{Y_{0}}}(L_{0},\Delta_{A},A,U_{\Delta_{Y_{0}}})$; $Y_{0}=f_{Y_{0}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},U_{Y_{0}})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},U_{L_{1}})$; $\Delta_{Y_{1}}=f_{\Delta_{Y_{1}}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},U_{\Delta_{Y_{1}}})$; $Y_{1}=f_{Y_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},\Delta_{Y_{1}},U_{Y_{1}})$.
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at follow-up under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at follow-up among those at risk at baseline: if $Y_{0}=0$, set $\Delta_{Y_{1}}=1$; else set $\Delta_{Y_{1}}=0$. Thereby, this final intervention is a dynamic or personalized one (e.g., [21, 34, 35, 36]). Denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects” and have been used in several real-data analyses [21, 5, 22, 23, 24, 25, 26]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [37, 38, 39, 40]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk.
To identify this effect, we re-express the conditional probability as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$ (1)
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$ with proof in eAppendix C.1. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at follow-up but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$):
  • the baseline and time-varying covariates capture the common causes of the joint outcome and follow-up measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e..
Under these assumptions and with proof given in eAppendix C.2, the numerator is identified as
$\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at follow-up is given
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
for exposure level $A=a$. Then we can define associations in terms of contrasts $\Psi(\mathbb{P};a)$ at different exposure levels. Concretely, in SEARCH-TB, we were interested in evaluating the association of alcohol consumption on incident TB infection with the risk ratio: $\Psi(\mathbb{P})=\Psi(\mathbb{P};1)\div\Psi(\mathbb{P};0)$.
3.5 Accounting for Participant Dependence
Finally, we outline an approach to account for dependence between participants within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, alcohol use may be influenced by members of a participant’s social circle, and TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [41, 42, 43, 44]). Importantly, by following the Causal Roadmap or a similar framework for causal inference [45, 46], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [47, 44, 22, 23, 48]).
Suppose it is reasonable to assume that participants are dependent within households, but households are effectively independent. (We relax this assumption below.) Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. In other settings, we may need to expand out the causal model and identification results to accommodate arbitrary dependence across a participant’s social network or across participants within a community. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [47, 23, 48].
4 Statistical Estimation and Inference
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as G-computation or inverse probability weighting (IPW) [12, 11]. Here, we highlight the use of targeted minimum loss-based estimation (TMLE), which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [14]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [49]. Thereby, TMLE harnesses machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [14]. When the $N$ participants are i.i.d., the estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [50]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand (i.e., numerator/denominator). Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [51]. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience. Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.
5 Application to SEARCH-TB
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets needed to account for confounding, missingness, and participant dependence [45, 52, 53, 31]. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by houshold [47, 23]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the impact of modeling assumptions, we also implemented the inverse probability weighting estimator (IPW) with the same adjustment set, but using parametric regressions to estimate the weights. Unlike TMLE, IPW is singly robust — relying on consistent estimation of the propensity score regressions — and tends to be inefficient [14]. Finally, to examine the impact of our missing data assumptions, we implemented the complete case analysis: the simple ratio of mean outcomes after restricting to participants whose exposure and outcomes were measured. For statistical inference, both IPW and the complete case approach accounted for clustering by household.
In the primary analysis (using TMLE with clustering by household), we found that alcohol use was associated with a 49% increase in the risk of incident TB, after flexibly accounting for confounding and missing data: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [31]. As shown in Figure 5, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62). In contrast, IPW relying on parametric assumptions resulted in a smaller association and very wide confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the complete case analysis was the least precise and resulted in very wide confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
Figure 5: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Complete Case” with an unadjusted estimator after excluding participants with missing exposures and outcomes.
6 Discussion
We presented causal models, causal parameters, and identification results for a series of observational studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding, missingness, and dependence, we found a 49% relative increase in the risk of incident TB associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). With the same adjustment approach but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, without adjustment, the complete case analysis yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
There are limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data application, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [54, 55, 56]. Future work is needed to investigate the assumptions, implementation, and performance of MI in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Third, we relied on various versions of the missing-at-random assumption throughout. In practice, data may be missing as a result of unobserved variables. When data are missing-not-at-random, we may need to collect additional data and conduct sensitivity/bias analyses [57, 10]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [45, 58, 52, 53]. Finally, we note that interpretation of causal estimands is nuanced when the exposure influences measurement and outcomes (Figures 3 and 4). By considering hypothetical interventions to ensure outcome measurement, we are blocking part of the exposure’s effect. In other words, our causal estimands correspond to a controlled direct effect — not the total effect. In a competing events setting, Young and colleagues have made similar points and defined causal estimands in terms of “separable” direct and indirect effects [15, 59, 60].
Overall, our goal was to contribute to the missing data literature by providing a framework to avoid complete case analyses and, instead, transparently state the missing data assumptions and robustly estimate the corresponding statistical estimands. To do so, we offered guidance for defining and identifying causal effects in real-world studies with missing exposures, missing outcomes, and dependence. We also demonstrated the real-world consequences of our causal and statistical assumptions in SEARCH-TB.
Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done. Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved. We also thank Dr. Diane Havlir and Dr. Maya L. Petersen, who together with Dr. Moses R. Kamya are the MPIs of the SEARCH collaboration.
7 Supplementary Digital Content for “Causal Inference with Missing Exposures, Missing Outcomes, and Dependence”
In the following, we provide proofs for the identification results. To match the applied example, we focus on binary outcomes, but our results generalize to all outcome-types. For simplicity we focus on categorical covariates, but our summations generalize to integrals for continuous covariates.
eAppendix A: Missing exposures (Figure 2 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp\Delta_{A}\mid L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp A\mid\Delta_{A}=1,L$
$=$ $\sum_{l}\mathbb{P}(Y=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$=$ $\mathbb{E}\big{[}\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)\big{]}$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
eAppendix B: Missing Exposures and Outcomes (Figure 3 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y^{*}\perp A\mid\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},\Delta_{A}=1,A=a,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid\Delta_{a}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\bigg{\{}\mathbb{E}\big{[}\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})\big{]}\bigg{\}}$
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$.
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption from the previous section.
eAppendix C: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 4 in the main text)
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$
Using the form of the target parameter on the right-hand side of the above equation, we proceed by presenting a separate identification result for the numerator and denominator separately.
C.1 Identification proof for the denominator
Under the following assumptions, which are analogous to eAppendix A, we can identify 1 minus the denominator:
$\mathbb{P}(Y_{0}^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y_{0}^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y_{0}^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Y_{0}^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{a}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\big{[}\mathbb{E}(Y\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$
along with the corresponding positivity assumptions.
C.2 Identification proof for the numerator
Let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$. Then under the following assumptions, we can identify the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)=\mathbb{P}(Z^{*}=1)$.
$\mathbb{P}(Z^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Z^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Z^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{y_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=y_{0},\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
by $\text{by }Z^{*}=0\text{ when }Y_{0}=1$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption for the denominator.

### v2

@@
-Causal Inference with Missing Exposures, Missing Outcomes, and Dependence
+Causal Inference with Missing Exposures and Missing Outcomes
 Abstract
 Missing data are ubiquitous in public health research.
-The missing-completely-at-random (MCAR) assumption is often unrealistic and can lead to meaningful bias when violated.
-The missing-at-random (MAR) assumption tends to be more reasonable, but guidance on conducting causal analyses under MAR is limited when there is missingness on multiple variables.
-We present a series of causal graphs and identification results to demonstrate the handling of missing exposures and outcomes in observational studies.
-For estimation and inference, we highlight the use of targeted minimum loss-based estimation (TMLE) with Super Learner to flexibly and robustly address confounding, missing data, and dependence.
+When estimating causal effects, there are well-established methods to address bias to due censored outcomes.
+Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent censoring.
+Identification is evaluated with the sequential backdoor criterion and considerations of data support.
+Then inverse weighting, standardization, and doubly-robust approaches are applied for statistical estimation and inference.
+We demonstrate how this framework can be extended to settings with missingness on the exposure of interest as well as the variable defining the population of interest (e.g., persons at risk of the outcome).
 Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda.
-This study posed notable challenges due to confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), missingness on the outcome at follow-up (capturing who acquired TB), and clustering within households.
-Application to real data from SEARCH-TB highlighted the real-world consequences of the discussed methods.
-Estimates from TMLE suggested that alcohol use was associated with a 49% increase in the relative risk (RR) of incident TB infection (RR=1.49, 95%CI: 1.39–1.59).
-These estimates were notably larger and more precise than estimates from inverse probability weighting (RR=1.13, 95%CI: 1.00–1.27) and unadjusted, complete case analyses (RR=1.18, 95%CI: 0.89–1.57).
-Our work demonstrates the utility of causal models for describing the missing data mechanism and TMLE for flexible inference.
-Keywords: Causal Inference, Missing Data, Targeted Minimum Loss-based Estimation (TMLE), Ensemble Machine Learning, Super Learner, Tuberculosis, Missing at Random
-1 Introduction
+This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), and missingness on the outcome at follow-up (capturing who acquired TB).
+We present a series of causal models and identification results to demonstrate the handling of missing exposures and outcomes in prospective studies.
+We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
+Keywords: Causal Inference, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
+Introduction
 Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8].
-A common approach is to simply exclude observations with missing data on the relevant variables.
-This “complete case” analysis can provide unbiased estimates if the data are missing as a result of a completely random process (i.e., missing-completely-at-random [MCAR]) [9].
-However, this assumption is often unrealistic in health settings, leading to biased estimates as well as reduced statistical power (e.g., [10]).
-The Missing-at-Random (MAR) assumption allows missingness to depend on observed study variables and is often more plausible [9].
-Suppose, for example, we aim to estimate the prevalence of a disease in the target population.
-On $N$ randomly sampled participants, we measure baseline covariates (e.g., age, gender, socio-economic status), but do not fully ascertain the outcome of interest.
-To recover the population-level outcome prevalence under MAR, we would need that the covariates capture all the common causes of measurement and outcomes as well as a positive probability of measurement within all possible covariate values.
-Then using standardization (a.k.a., G-computation), inverse weighting, multiple imputation, or more robust methods, we could obtain a point estimate and 95% confidence intervals using data on all $N$ participants.
-Instead, if we were willing to consider the stronger MCAR assumption, corresponding to no common causes of measurement and the outcome, then we could estimate disease prevalence with an empirical proportion among those measured.
-This prevalence example illustrates the consequences of causal assumptions on missing data: stronger assumptions allow for simpler estimation approaches, while weaker assumptions require more complex estimation approaches.
-An analogous scenario arises in analyses of observational data, which are subject to confounding.
-In the following, we present a comprehensive framework to address common real-world challenges: confounding, missingness on multiple variables, and dependence between participants in our study.
-Our work builds on a long history of methods to address missing data (e.g., [11, 9, 12, 13, 14, 15]).
-As detailed below, we give special consideration to missing exposures.
-To the best of our knowledge, methods to handle missing exposure have largely focused on case-control designs, where exposure information is retrospectively collected on a sample of participants based on their outcome status (e.g., [16, 17, 4, 18, 19, 20]).
-Here, we focus on standard cross-sectional or longitudinal studies with missing exposures.
+There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]).
+The Causal Roadmap provides one such approach [21, 22, 23].
+First, we specify our causal question: how would expected outcomes differ if all did versus did not receive the exposure of interest and censoring were prevented.
+Then, we specify a causal model, such as directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for observed data: baseline and time-varying confounders, exposures, and outcomes [10].
+Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement.
+Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold) [9, 10, 11, 12].
+Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient).
+Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly-robust methods, such as targeted minimum loss-based estimation (TMLE) [11].
+Finally, we conduct sensitivity analyses and appropriately interpret the results.
+Here, we provide a tutorial on how this framework can be extended to address missingness on the exposure.
 We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest.
 Suppose, for example, we are interested in studying the incidence of some disease.
-Our target population would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline.
-In this setting, our incidence estimates would be subject to bias if there is differential measurement of outcomes at baseline.
+Our population of interest would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline.
+In this setting, our incidence estimates would be subject to bias if there is differential missingness of outcomes at baseline.
 Using Counterfactual Strata Effects [24, 5, 25, 26, 27, 28, 29], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios.
-Finally, we provide extensions for settings where participants are not independent and, instead, clustered within households or communities.
-Altogether, our presentation covers multi-source missing data, confounding, and dependence.
-We build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers.
-We illustrate the practical relevance with the SEARCH-TB study.
-2 Motivating example
+We note there is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”) [30, 31].
+In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link [4].
+They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures [4, 32].
+Holovchak et al.
+extended this work for the effects of longitudinal exposures [33], while questioning the practical utility of the m-DAG approach.
+Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis” [33].
+As an alternative, we demonstrate how an established framework for causal inference can used to define, identify, and estimate causal estimands, such as the average treatment effect, under missingness on exposures and outcomes.
+In the following, we build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers who need to address confounding and multi-source missingness.
+We illustrate practical relevance with the SEARCH-TB study.
+Motivating example
 SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [34].
@@
 Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection.
-Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection, a proxy for population-level transmission, was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [29, 30].
+Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [36, 37].
 This sub-study was intentionally enriched for persons with HIV.
@@
 One year later, follow-up tests were administered to participants who tested negative at baseline.
-Here, we provide an in-depth exploration of the methods to evaluate the effect of alcohol use on incident TB infection in SEARCH-TB [31]: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population of interest, (3) missingness in the final outcome, and (4) confounding.
-Additionally, given our focus on infectious disease outcomes, the independent and identically distributed (i.i.d.) assumption was likely violated, and we accounted dependence among study participants.
-3 Related Causal Problems: Building Complexity
+Here, we demonstrate the methods used to evaluate the effect of alcohol use on incident TB infection [38]: confounding, missingness on the exposure of interest, missingness on the baseline outcome (defining who was at risk of TB), and missingness on the final outcome (defining who acquired TB).
+Related Causal Problems: Building Complexity
 Many studies feature only a subset of the challenges described above.
 We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies.
-For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or the causal risk ratio).
-3.1 Classic point-treatment problem
-First, we consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close.
+For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
+Classic point-treatment problem
+In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close.
 This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline.
-The directed acylic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 1 and reflect the simplifying assumption that participants are independent (e.g., cannot transmit TB to one another due to geographical distance).
-Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest.
-Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$.
-Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$.
-In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
-Figure 1: Directed acyclic graph (DAG) for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$.
-The corresponding non-parametric structural equation model (NPSEM) is given by $L=f_{L}(U_{L})$; $A=f_{A}(L,U_{A})$; $Y=f_{Y}(L,A,U_{Y})$ where $(U_{L},U_{A},U_{Y})$ represent unmeasured factors determining the values of the covariates, exposure, and outcome, respectively.
-On the DAG, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$.
-To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$ and can be represented as $Y^{a}\perp A\mid L$.
-Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e..
-Under these two assumptions, our causal target is equal to the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [12].
-Even if these assumptions are not reasonable (e.g., there are unmeasured confounders in Figure 1), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
-3.2 Missing exposures
+Let $Y^{a}$ be the counterfactual outcome if, possibly contrary-to-fact, the participant had exposure-level $A=a$.
+The counterfactual mean outcome $\mathbb{E}(Y^{a})$ is identified if there are no unmeasured confounders and sufficient data support: $Y^{a}\perp A\mid L$ and $\mathbb{P}(A=a\mid L)>0$, respectively.
+The statistical estimand is given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$[9].
+Even if these assumptions are not reasonable, the G-computation formula is still a well-defined statistical estimand, on which we can focus our estimation efforts (Appendix S1).
+Missing Outcomes
 Most real-world studies, however, depart from this idealized point-treatment problem.
-The first departure we consider is missingness on the exposure of interest.
+Here, we review causal models and identification when there is missingness on the outcome.
+Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not test for TB at the end of follow-up.
+Let $\Delta_{Y}$ be an indicator of outcome measurement.
+If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual.
+However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed.
+In Figure 1, we show two possible ways of using causal models to represent such a study.
+In Panel A, we introduce $Y^{\circ}$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times Y^{\circ}$.
+Thus, $Y$ equals its underlying value $Y^{\circ}$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$).
+In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$.
+In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because when an outcome is not measured, it is set deterministically to NA.
+Both approaches lead us to same identification assumptions and statistical estimands.
+Therefore, to minimize notation and mirror the long-standing literature on censoring (e.g., [16, 17, 11, 18, 19, 20, 12]), we use the latter representation in the remainder of the article.
+To define causal effects when the outcome is subject to missingness, we now consider counterfactuals indexed by both the exposure and the outcome measurement indicator.
+Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented.
+Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had exposure $A=a$ and their outcomes measured.
+To identify this causal parameter and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness.
+Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement.
+These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively.
+These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion [9, 10, 11, 12].
+We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e.
+and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively.
+If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$ [9].
+Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $Y^{\circ}$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, YY= observed outcome.
+For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
+Missing Exposures
+We now demonstrate how the above approach for missing outcomes can be applied to missing exposures.
 Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use.
@@
 If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual.
-However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is not observed (i.e., equal to NA).
-The DAG and NPSEM for such a study are given in Figure 2.
-To define causal effects when the exposure is subject to missingness, we now consider counterfactuals indexed by both the exposure and its measurement indicator.
-Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and was at level $A=a$.
-Figure 2: DAG for a point-treatment problem with missingness on the exposure: $L$=baseline covariates, $\Delta_{A}$=indicator of exposure measurment, $A$=exposure, $Y$=outcome, and $U^{*}$=unmeasured common causes of at least two variables in $(L,\Delta_{A},A,Y)$.
-The corresponding NPSEM is $L=f_{L}(U_{L})$; $\Delta_{A}=f_{\Delta_{A}}(L,U_{\Delta_{A}})$, $A=f_{A}(L,\Delta_{A},U_{A})$, $Y=f_{Y}(L,\Delta_{A},A,U_{Y})$.
-Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had their exposure measured $\Delta_{A}=1$ and it were at level $A=a$.
-To identify this causal parameter and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we would need $L$ to be sufficient to control for confounding and differential measurement.
-In more detail, the following assumptions are required:
-  • Within values of the covariates $L$, outcomes among those with measured exposures are representative of outcomes among those without measured exposures: $Y^{*}\perp\Delta_{A}\mid L$.
-  • There is a non-zero probability of exposure measurement within all possible values of $L$: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e..
-  • Among those with measured exposures ($\Delta_{A}=1$),
-    • $L$ captures all the common causes of the exposure and outcome: $Y^{*}\perp A\mid\Delta_{A}=1,L$.
-    • there is a non-zero probability of being exposed to level $A=a$ within all possible values of $L$: $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e..
-If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in eAppendix A.
-As before, even if these assumptions do not hold (e.g., there are unmeasured common causes of measurement and the outcome in Figure 2), we still have a well-defined statistical estimand for estimation and inference.
-3.3 Missing Exposures and Outcomes
-We now add another common complication: missing outcomes.
+However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is set deterministically to NA.
+The DAG and NPSEM for such a study are given in Figure 1C.
+To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator.
+Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
+Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection on missing outcomes.
+Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively.
+These assumptions can be graphically evaluated with the sequential backdoor criterion.
+We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e.
+and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e.
+If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S2.1.
+Missing Exposures and Outcomes
+We now combine our two challenges: missing exposures and missing outcomes.
 Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment.
 To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting.
-Let $L_{0}$ be baseline covariates, $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment, and $\Delta_{Y}$ be an indicator of outcome measurement.
-Specifically, if $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as before.
-However, if $\Delta_{Y}=0$ for a participant, their outcome $Y$ is not observed (i.e., equal to NA).
-The simplified DAG and full NPSEM for such a study are given in Figure 3.
-Figure 3: DAG with missingness on the exposure and outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, $A$=exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of measured outcome, and $Y$=outcome.
-For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least two variables in $(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$.
-The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,U_{L_{1}})$; $\Delta_{Y}=f_{\Delta_{Y}}(L_{0},\Delta_{A},A,L_{1},U_{\Delta_{Y}})$; $Y=f_{Y}(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},U_{Y})$.
+Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment.
+Figure 2 provides the causal models for such a study.
+Figure 2: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of outcome measurement, and YY=observed outcome.
+For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
 To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators.
@@
 Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship.
-Therefore, we rely on sequential randomization and find a set of covariates that satisfies the backdoor criteria for each “intervention” node given the observed past [12].
+Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past [9].
 As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding.
 In other words, we need the analogous identification assumptions given in the prior subsection.
-Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$),
-  • the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$.
-  • there is a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e..
-If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in eAppendix B [12, 32, 33].
-As before, even if these identification assumptions do not hold, we still have a well-defined statistical estimand to focus our estimation efforts.
-3.4 Missing Exposures and Outcomes at Baseline and Follow-up
-We now consider missingness on the outcome at baseline.
-Continuing our running example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing.
-In other words, our cohort of participants without TB at baseline is subject to selection bias.
+Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively.
+If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S2.2 [9, 39, 40].
+Missing Exposures and Missing Outcomes at the Start and End of Follow-up
+We now consider missingness on the outcome at the start of follow-up.
+Following our motivating example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing.
+In other words, our longitudinal cohort of participants without TB is subject to selection bias.
 To reflect this data generating process, we update our notation to have multiple outcome measures.
-Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at baseline, $Y_{0}$ be an indicator having the outcome at baseline, $\Delta_{Y_{1}}$ be an indicator of outcome measurement at follow-up, and $Y_{1}$ be an indicator having the outcome at follow-up.
-As previously introduced, the value of the baseline outcome defines the target population; specifically, we are interested in occurrence of the outcome ($Y_{1}=1$) among those who are at risk at baseline ($Y_{0}=0$).
+Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”).
+Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes; they are set to NA when not observed.
 The corresponding causal models can be found in Figure 3.
-Figure 4: DAG with missingness on the exposure, the baseline outcome, and the follow-up outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, A=exposure, $\Delta_{Y_{0}}$=indicator of measured baseline outcome, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of measured outcome at follow-up, and $Y_{1}$= outcome at follow-up.
-For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least 2 variables in $(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$.
-The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $\Delta_{Y_{0}}=f_{\Delta_{Y_{0}}}(L_{0},\Delta_{A},A,U_{\Delta_{Y_{0}}})$; $Y_{0}=f_{Y_{0}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},U_{Y_{0}})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},U_{L_{1}})$; $\Delta_{Y_{1}}=f_{\Delta_{Y_{1}}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},U_{\Delta_{Y_{1}}})$; $Y_{1}=f_{Y_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},\Delta_{Y_{1}},U_{Y_{1}})$.
+The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$.
+As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at risk of TB at baseline ($Y_{0}=0$).
+Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline.
+For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
 To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$.
-Additionally, we consider the counterfactual outcome at follow-up under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at follow-up among those at risk at baseline: if $Y_{0}=0$, set $\Delta_{Y_{1}}=1$; else set $\Delta_{Y_{1}}=0$.
-Thereby, this final intervention is a dynamic or personalized one (e.g., [21, 34, 35, 36]).
-Denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
+Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at risk at baseline.
+Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline.
+Thereby, this final intervention is a dynamic or personalized one to set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise (e.g., [24, 41, 42, 43]).
+For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
 Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$.
-Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects” and have been used in several real-data analyses [21, 5, 22, 23, 24, 25, 26].
+Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”[24, 5, 25, 26, 27, 28], which are defined by Nakato et al.
+as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure” [29].
 These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [44, 45, 46, 47].
 For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use.
-Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk.
+Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk at baseline.
 To identify this effect, we re-express the conditional probability as
@@
 The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk.
-The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the outcome.
-Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$ with proof in eAppendix C.1.
+The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome.
+Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S2.3.
 Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
 In our final causal parameter (Eq.
-1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at follow-up but not at baseline.
+1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at endline but not at baseline.
 For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values.
 To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following.
-Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$):
-  • the baseline and time-varying covariates capture the common causes of the joint outcome and follow-up measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
-  • there is a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e..
-Under these assumptions and with proof given in eAppendix C.2, the numerator is identified as
+Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
+We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e.
+Under these assumptions and with proof given in Appendix S2.3, the numerator is identified as
 $\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
-Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at follow-up is given
+Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at endline is given by
 $\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
@@
 Concretely, in SEARCH-TB, we were interested in evaluating the association of alcohol consumption on incident TB infection with the risk ratio: $\Psi(\mathbb{P})=\Psi(\mathbb{P};1)\div\Psi(\mathbb{P};0)$.
-3.5 Accounting for Participant Dependence
-Finally, we outline an approach to account for dependence between participants within groups or clusters, such as households, schools, hospitals, or communities.
-Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases.
-In our running example, alcohol use may be influenced by members of a participant’s social circle, and TB is transmitted from person to person.
-This dependence should be reflected in the corresponding causal model (e.g., [41, 42, 43, 44]).
-Importantly, by following the Causal Roadmap or a similar framework for causal inference [45, 46], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [47, 44, 22, 23, 48]).
-Suppose it is reasonable to assume that participants are dependent within households, but households are effectively independent.
-(We relax this assumption below.) Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors.
-Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$.
-In other settings, we may need to expand out the causal model and identification results to accommodate arbitrary dependence across a participant’s social network or across participants within a community.
-The exact form of the causal model and identification result will depend on the application.
-Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [47, 23, 48].
-4 Statistical Estimation and Inference
+Statistical Estimation and Inference
 In the previous section, we introduced a series of causal models and identification results of increasing complexity.
-For the resulting statistical estimands, we could use a singly robust estimation approach, such as G-computation or inverse probability weighting (IPW) [12, 11].
-Here, we highlight the use of targeted minimum loss-based estimation (TMLE), which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [14].
+For the resulting statistical estimands, we could use a singly robust estimation approach, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW) [9, 13].
+Here, we highlight the use of TMLE, which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [11].
 In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation.
 Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [48].
-Thereby, TMLE harnesses machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions.
+Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions.
 Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq.
@@
 TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [11].
-When the $N$ participants are i.i.d., the estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [50].
+The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [49].
 The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values.
@@
 For ratio-type estimands (Eq.
-2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand (i.e., numerator/denominator).
+2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand.
 Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
-If clustering is present, estimation and inference must be adjusted.
-First, the cross-validation scheme used within Super Learner must respect the independent unit.
-Concretely, participants in a given cluster are all assigned to the same sample-split.
-Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [51].
-Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.
-Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$.
-We then proceed with variance estimation using the cluster-level influence curve.
-This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.
-5 Application to SEARCH-TB
+Application to SEARCH-TB
 We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda?
-With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets needed to account for confounding, missingness, and participant dependence [45, 52, 53, 31].
+With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets [21, 22, 23, 50, 38].
+Concretely, the causal model reflected team’s knowledge of study protocol, the study procedures, and the epidemiology of TB in the region.
 Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures.
 For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean.
-We conducted influence curve-based inference, accounting for clustering by houshold [47, 23].
+We conducted influence curve-based inference, accounting for clustering by household (Appendix S4) [51, 26].
 In secondary analyses, we considered communities, instead of households, to be the independent unit.
-To examine the impact of modeling assumptions, we also implemented the inverse probability weighting estimator (IPW) with the same adjustment set, but using parametric regressions to estimate the weights.
-Unlike TMLE, IPW is singly robust — relying on consistent estimation of the propensity score regressions — and tends to be inefficient [14].
-Finally, to examine the impact of our missing data assumptions, we implemented the complete case analysis: the simple ratio of mean outcomes after restricting to participants whose exposure and outcomes were measured.
-For statistical inference, both IPW and the complete case approach accounted for clustering by household.
-In the primary analysis (using TMLE with clustering by household), we found that alcohol use was associated with a 49% increase in the risk of incident TB, after flexibly accounting for confounding and missing data: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [31].
-As shown in Figure 5, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62).
-In contrast, IPW relying on parametric assumptions resulted in a smaller association and very wide confidence intervals: RR=1.13 (95%CI: 1.00-1.27).
-Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the complete case analysis was the least precise and resulted in very wide confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
-Figure 5: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Complete Case” with an unadjusted estimator after excluding participants with missing exposures and outcomes.
-6 Discussion
-We presented causal models, causal parameters, and identification results for a series of observational studies with increasing levels of missingness.
+To examine the sensitivity of our results to alternative estimation approaches, we also implemented IPW for the same statistical estimand, but using parametric regressions to estimate the weights.
+Finally, to examine the impact of handling of our missing data, we took the following “naïve” approach: (1) subset on participants known to be at risk of TB at baseline ($Y_{0}=0$); (2) further subset on participants with measured exposures ($\Delta_{A}=1$) and measured outcomes at endline ($\Delta_{Y_{1}}=1$), and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)\mid Y_{0}=0]$.
+This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3).
+Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding [52, 53, 54].
+For statistical inference, both IPW and the naïve approach accounted for clustering by household.
+In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [38].
+As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S4.
+In contrast, IPW relying on parametric assumptions resulted in a smaller association and confidence intervals overlapping the null: RR=1.13 (95%CI: 1.00-1.27).
+Finally, after restricting to participants who tested negative at baseline, responded to questions about alcohol use, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
+Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at risk at baseline and with measured exposures and outcomes at endline.
+Discussion
+We presented causal models, causal parameters, and identification results for a series of prospective studies with increasing levels of missingness.
 For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands.
 Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work.
-Using TMLE to flexibly account for confounding, missingness, and dependence, we found a 49% relative increase in the risk of incident TB associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59).
-With the same adjustment approach but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27).
-Finally, without adjustment, the complete case analysis yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
-There are limitations to our work.
+Using TMLE to flexibly account for confounding and missingness, we found a 49% relative increase in the risk of incident TB infection associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59).
+For the same statistical estimand but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27).
+Finally, the naïve approach to condition on the participants with baseline TB risk and measured exposures/outcomes yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
+There are several advantages to our framework for handling missing exposures and outcomes when evaluating the causal effects.
+Unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data [4], our approach focuses our efforts on the portions of the observed data distribution relevant for the statistical estimand.
+Furthermore, our approach uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (sometimes called an “available-case analysis”) [4].
+Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
+A major limitation to this work is that we did not address missingness on the confounders.
+Some approaches include mean/median imputation or including missingness indicators in the adjustment set.
+For further discussion on missing confounders, we refer to the comprehensive work of Williamson et al.
+[55].
+We also focused on cross-sectional or prospective studies.
+Thus, we did not cover scenarios where the outcome impacts the measurement of other variables.
+Such scenarios would arise in case-control studies and have been addressed in prior literature (e.g., [56, 57, 58, 59, 54]).
+There are other limitations to our work.
 First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health.
-Second, in our real-data application, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [54, 55, 56].
-Future work is needed to investigate the assumptions, implementation, and performance of MI in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants.
-Third, we relied on various versions of the missing-at-random assumption throughout.
-In practice, data may be missing as a result of unobserved variables.
-When data are missing-not-at-random, we may need to collect additional data and conduct sensitivity/bias analyses [57, 10].
+Second, in our real-data demonstration, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [60, 61, 62].
+Future work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants.
+Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness.
+In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses [63, 64].
 Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [21, 65, 23, 50].
-Finally, we note that interpretation of causal estimands is nuanced when the exposure influences measurement and outcomes (Figures 3 and 4).
-By considering hypothetical interventions to ensure outcome measurement, we are blocking part of the exposure’s effect.
-In other words, our causal estimands correspond to a controlled direct effect — not the total effect.
-In a competing events setting, Young and colleagues have made similar points and defined causal estimands in terms of “separable” direct and indirect effects [15, 59, 60].
-Overall, our goal was to contribute to the missing data literature by providing a framework to avoid complete case analyses and, instead, transparently state the missing data assumptions and robustly estimate the corresponding statistical estimands.
-To do so, we offered guidance for defining and identifying causal effects in real-world studies with missing exposures, missing outcomes, and dependence.
-We also demonstrated the real-world consequences of our causal and statistical assumptions in SEARCH-TB.
+Overall, our goal was to demonstrate how an established framework for estimating causal effects with censored outcomes could be extended to settings with missingness on the exposure, the outcome at baseline, and the outcome at endline.
+To do so, we provided a series of causal models with increasing complexity and discussed the identification assumptions for each.
+For estimation and inference, we focused on TMLE, a doubly-robust approach using machine learning and data from all participants (including those with missing exposures or outcomes).
+Our work was motivated by the investigation of the effect of alcohol use on incident TB infection [38], and our re-analysis demonstrated the real-world consequences of our approach.
 Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
-Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done.
-Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
+Conflicts of Interest: The authors report no conflict of interest in this work.
 Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved.
@@
 Kamya are the MPIs of the SEARCH collaboration.
-7 Supplementary Digital Content for “Causal Inference with Missing Exposures, Missing Outcomes, and Dependence”
+Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done.
+Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
+Supplementary Materials for “Causal Inference with Missing Exposures and Missing Outcomes”
+Contents:
+  • Appendix S1: More on the classic point-treatment problem
+  • Appendix S2: Proofs
+  • Appendix S3: Accounting for Outcome Dependence
+Appendix S1: More on the classic point-treatment problem
+We consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close.
+This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline.
+The directed acyclic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 5A.
+Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest.
+Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$.
+Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$.
+In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
+Figure 5: Causal models for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$.
+On the directed acyclic graph, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$.
+In panel A provides the causal models under no assumptions about the unmeasured factors.
+Panels B and C are compatible with the no unmeasured confounders assumption.
+To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$.
+This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$ [9, 10, 11, 12].
+Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$.
+In Figure 5B and C, we show two causal models where this assumption would hold.
+Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e.
+This is a condition on data support and known as the “positivity assumption”.
+We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model.
+Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [9].
+Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
+In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation [21, 22, 23, 66, 65] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
+Appendix S2: Proofs
 In the following, we provide proofs for the identification results.
@@
 For simplicity we focus on categorical covariates, but our summations generalize to integrals for continuous covariates.
-eAppendix A: Missing exposures (Figure 2 in the main text)
+Appendix S2.1: Missing exposures (Figure 1C in the main text)
 Let $Y^{*}=Y^{\Delta_{A}=1,A=a}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
@@
 $\text{by }Y^{*}\perp A\mid\Delta_{A}=1,L$
+$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
+by the consistency assumption
 $=$ $\sum_{l}\mathbb{P}(Y=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
 $=$ $\mathbb{E}\big[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)\big]$
-For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
-eAppendix B: Missing Exposures and Outcomes (Figure 3 in the main text)
+We again note the consistency assumption holds by our definition of counterfactual outcomes as being derived through interventions on the causal model.
+For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e.
+and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
+Appendix S2.2: Missing Exposures and Outcomes (Figure 2 in the main text)
 Let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$.
@@
 For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e.
-in addition to the positivity assumption from the previous section.
-eAppendix C: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 4 in the main text)
+in addition to the positivity assumptions from the previous section.
+Appendix S2.3: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 3 in the main text)
 Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as
@@
 Using the form of the target parameter on the right-hand side of the above equation, we proceed by presenting a separate identification result for the numerator and denominator separately.
-C.1 Identification proof for the denominator
-Under the following assumptions, which are analogous to eAppendix A, we can identify 1 minus the denominator:
+Identification proof for the denominator
+Under the following assumptions, which are analogous to Appendix S2.1, we can identify 1 minus the denominator:
 $\mathbb{P}(Y_{0}^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
@@
 along with the corresponding positivity assumptions.
-C.2 Identification proof for the numerator
+Identification proof for the numerator
 Let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$. Then under the following assumptions, we can identify the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)=\mathbb{P}(Z^{*}=1)$.
@@
 $\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
-by $\text{by }Z^{*}=0\text{ when }Y_{0}=1$
+$\text{by }Z^{*}=0\text{ when }Y_{0}=1$
 $=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
@@
 For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e.
-in addition to the positivity assumption for the denominator.
+in addition to the positivity assumptions for the denominator.
+Appendix S3: Accounting for Outcome Dependence
+Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities.
+Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases.
+In our running example, TB is transmitted from person to person.
+This dependence should be reflected in the corresponding causal model (e.g., [67, 68, 69, 70]).
+By following the Causal Roadmap or a similar framework for causal inference [21, 71], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [51, 70, 25, 26, 72]).
+Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households.
+Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors.
+(See, for example, [70].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$.
+The exact form of the causal model and identification result will depend on the application.
+Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [51, 26, 72].
+If clustering is present, estimation and inference must be adjusted.
+First, the cross-validation scheme used within Super Learner must respect the independent unit.
+Concretely, participants in a given cluster are all assigned to the same sample-split.
+Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [18].
+Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.
+Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$.
+We then proceed with variance estimation using the cluster-level influence curve.
+This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.

### v3

@@
 Missing data are ubiquitous in public health research.
-When estimating causal effects, there are well-established methods to address bias to due censored outcomes.
-Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent censoring.
-Identification is evaluated with the sequential backdoor criterion and considerations of data support.
-Then inverse weighting, standardization, and doubly-robust approaches are applied for statistical estimation and inference.
-We demonstrate how this framework can be extended to settings with missingness on the exposure of interest as well as the variable defining the population of interest (e.g., persons at risk of the outcome).
+When estimating causal effects, there are well-established methods to address bias to due missing outcomes.
+Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent missingness.
+We demonstrate how this framework can be extended to missing exposures.
+We further extend this framework to incorporate missingness on the baseline outcome, which induces missingness on the population of interest.
+To do so, we highlight the use of Counterfactual Strata Effects: causal estimands where the focus population is subject to missingness and/or impacted by the exposure.
 Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda.
-This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), and missingness on the outcome at follow-up (capturing who acquired TB).
-We present a series of causal models and identification results to demonstrate the handling of missing exposures and outcomes in prospective studies.
+This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at-risk of TB and, thus, in the focus population), and missingness on the outcome at follow-up (capturing who acquired TB).
+We present a series of causal models and identification results to demonstrate the handling of missingness in these settings.
 We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
-Keywords: Causal Inference, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
+Keywords: Causal Inference, Counterfactual Strata Effects, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
 Introduction
-Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8].
-There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]).
-The Causal Roadmap provides one such approach [21, 22, 23].
-First, we specify our causal question: how would expected outcomes differ if all did versus did not receive the exposure of interest and censoring were prevented.
-Then, we specify a causal model, such as directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for observed data: baseline and time-varying confounders, exposures, and outcomes [10].
-Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement.
-Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold) [9, 10, 11, 12].
-Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient).
-Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly-robust methods, such as targeted minimum loss-based estimation (TMLE) [11].
-Finally, we conduct sensitivity analyses and appropriately interpret the results.
-Here, we provide a tutorial on how this framework can be extended to address missingness on the exposure.
-We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest.
-Suppose, for example, we are interested in studying the incidence of some disease.
-Our population of interest would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline.
-In this setting, our incidence estimates would be subject to bias if there is differential missingness of outcomes at baseline.
-Using Counterfactual Strata Effects [24, 5, 25, 26, 27, 28, 29], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios.
-We note there is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”) [30, 31].
-In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link [4].
-They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures [4, 32].
-Holovchak et al.
-extended this work for the effects of longitudinal exposures [33], while questioning the practical utility of the m-DAG approach.
-Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis” [33].
-As an alternative, we demonstrate how an established framework for causal inference can used to define, identify, and estimate causal estimands, such as the average treatment effect, under missingness on exposures and outcomes.
-In the following, we build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers who need to address confounding and multi-source missingness.
-We illustrate practical relevance with the SEARCH-TB study.
-Motivating example
-SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [34].
-Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [35].
-Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection.
-Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [36, 37].
-This sub-study was intentionally enriched for persons with HIV.
-Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV.
-At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households.
-One year later, follow-up tests were administered to participants who tested negative at baseline.
-Here, we demonstrate the methods used to evaluate the effect of alcohol use on incident TB infection [38]: confounding, missingness on the exposure of interest, missingness on the baseline outcome (defining who was at risk of TB), and missingness on the final outcome (defining who acquired TB).
-Related Causal Problems: Building Complexity
+Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products.[37, 80, 69, 46, 5, 12, 33, 42] There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes.[61, 52, 76, 28, 31, 67, 73, 59, 70, 68, 77, 84] The Causal Roadmap provides one such approach.[53, 15, 14, 21] The standard implementation of the Roadmap with censored outcomes is as follows.
+First, we specify our causal question: how would expected outcomes differ if all did versus did not have the exposure of interest and censoring were prevented.
+Then, we specify a causal model, such as a directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for the baseline and time-varying confounders, exposures, and outcomes.[52, 76] Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement.
+Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold).[61, 52, 76, 28] Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient).
+Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly robust methods, such as targeted minimum loss-based estimation (TMLE).[76] Finally, we conduct sensitivity analyses and appropriately interpret the results.
+Here, we first discuss how this framework for missing and censored outcomes can be extended to address missingness on the exposure.
+We then introduce Counterfactual Strata Effects: causal estimands defined for a group which is subject to missingness and/or influenced by the exposure.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] Specifically, we highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest.
+Concretely, suppose we are interested in studying the incidence of some disease.
+Our focus population would be persons who are at-risk of the developing the outcome and are, thereby, disease-free at baseline.
+In this setting, our incidence estimates would be subject to bias if our analyses did not account for differential missingness of outcomes at baseline.
+We provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios—overall and, importantly, when the exposure influences the baseline risk of the outcome.
+We also discuss how our approach avoids “failures” sometimes arising in the target trial emulation framework.[26]
+We illustrate practical relevance with SEARCH-TB, which aimed to evaluate the effect of alcohol use on incident tuberculosis (TB) infection.[1] SEARCH-TB was a sub-study of SEARCH, a cluster randomized trial to evaluate a community-based approach to Universal HIV Test-and-Treat in rural Kenya and Uganda (2013-2017; NCT01864603).[25] At baseline in SEARCH, we conducted a rapid census and then population-based measurement of sociodemographic factors (e.g., age, sex, mobility, and alcohol use) and HIV.[11] Due to high costs and complex logistics, evaluation of incident TB infection was limited to eastern Ugandan and intentionally enriched for persons with HIV.[39, 40] Specifically, we over-sampled households with at least one adult with HIV.
+Then we aimed to administer tuberculin skin tests to residents of the sampled households.
+One year later, we aimed to administer follow-up tests to participants who tested negative at baseline in the sub-study.
+Here, we demonstrate the methods used to address several real-world challenges arising in SEARCH-TB: confounding, missingness on the exposure of interest (self-report of any alcohol use), missingness on the baseline outcome (defining who was at-risk of TB), and missingness on the final outcome (defining who acquired TB).
+Building to Counterfactual Strata Effects
 Many studies feature only a subset of the challenges described above.
-We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies.
-For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
-Classic point-treatment problem
+We provide causal models and identification results for a series of hypothetical studies with increasing complexity.
+For simplicity, we focus on defining and identifying causal estimands under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
 In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close.
-This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline.
-Let $Y^{a}$ be the counterfactual outcome if, possibly contrary-to-fact, the participant had exposure-level $A=a$.
-The counterfactual mean outcome $\mathbb{E}(Y^{a})$ is identified if there are no unmeasured confounders and sufficient data support: $Y^{a}\perp A\mid L$ and $\mathbb{P}(A=a\mid L)>0$, respectively.
-The statistical estimand is given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$[9].
-Even if these assumptions are not reasonable, the G-computation formula is still a well-defined statistical estimand, on which we can focus our estimation efforts (Appendix S1).
+Appendix S2 provides a brief review of approaches to handle missingness on the baseline confounders $L$.
 Missing Outcomes
-Most real-world studies, however, depart from this idealized point-treatment problem.
-Here, we review causal models and identification when there is missingness on the outcome.
-Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not test for TB at the end of follow-up.
+Consider our motivating study of the effect of alcohol use on incident TB infection.
+Suppose that we have a representative cohort of persons without TB at baseline, but some participants did not test for TB at the end of follow-up.
 Let $\Delta_{Y}$ be an indicator of outcome measurement.
@@
 However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed.
-In Figure 1, we show two possible ways of using causal models to represent such a study.
-In Panel A, we introduce $Y^{\circ}$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times Y^{\circ}$.
-Thus, $Y$ equals its underlying value $Y^{\circ}$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$).
+Figure 1 provides two possible causal models to represent such a study.
+In Panel A, we use $\Upsilon$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times\Upsilon$.
+Thus, $Y$ equals its underlying value $\Upsilon$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$).
 In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$.
-In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because when an outcome is not measured, it is set deterministically to NA.
+In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because the outcome is missing if $\Delta_{Y}=0$.
 Both approaches lead us to same identification assumptions and statistical estimands.
-Therefore, to minimize notation and mirror the long-standing literature on censoring (e.g., [16, 17, 11, 18, 19, 20, 12]), we use the latter representation in the remainder of the article.
-To define causal effects when the outcome is subject to missingness, we now consider counterfactuals indexed by both the exposure and the outcome measurement indicator.
+Therefore, to minimize notation and mirror the long-standing literature on censoring, we use the latter representation in the remainder of the article.[60, 59, 8, 70, 76, 68, 77, 84, 28]
+To define causal effects when the outcome is subject to missingness, we consider counterfactuals indexed by both the exposure and the outcome measurement indicator.
 Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented.
-Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had exposure $A=a$ and their outcomes measured.
-To identify this causal parameter and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness.
+Then our causal estimand $\mathbb{E}(Y^{*})$ is the expected counterfactual outcome if all participants had exposure $A=a$ and their outcomes measured.
+To identify this causal estimand and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness.
 Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement.
 These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively.
-These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion [9, 10, 11, 12].
-We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e.
+These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion.[61, 52, 76, 28] We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e.
 and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively.
-If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$ [9].
-Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $Y^{\circ}$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, YY= observed outcome.
+If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$.[61]
+Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $\Upsilon$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, and YY= observed outcome.
 For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
 If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual.
-However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is set deterministically to NA.
+However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is missing.
 The DAG and NPSEM for such a study are given in Figure 1C.
@@
 Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
-Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection on missing outcomes.
+Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection.
 Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively.
-These assumptions can be graphically evaluated with the sequential backdoor criterion.
 We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e.
@@
 We now combine our two challenges: missing exposures and missing outcomes.
-Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment.
-To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting.
+Continuing our running example, suppose that among our cohort of persons at-risk of TB, some participants did not answer questions about their alcohol use and some participants did not test for TB at the end of follow-up.
+To reflect this data generating process, we introduce new notation to reflect the longitudinal setting.
 Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment.
@@
 Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement.
-To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders.
+To identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders.
 Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship.
-Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past [9].
-As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding.
+Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past.[61] As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding.
 In other words, we need the analogous identification assumptions given in the prior subsection.
 Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively.
-If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S2.2 [9, 39, 40].
-Missing Exposures and Missing Outcomes at the Start and End of Follow-up
-We now consider missingness on the outcome at the start of follow-up.
-Following our motivating example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing.
-In other words, our longitudinal cohort of participants without TB is subject to selection bias.
-To reflect this data generating process, we update our notation to have multiple outcome measures.
+If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S3.2.[61, 8, 72]
+Counterfactual Strata Effects
+Following our motivating example, we now introduce missingness on the outcome at baseline, which induces missingness on the focus population.
+Specifically in SEARCH-TB, we were interested in the effect of alcohol use among all persons at-risk of TB, but we did not reach all for initial TB testing.
+Therefore, our longitudinal cohort of participants who tested TB-negative at baseline in the sub-study was not representative of all persons at-risk of TB.
+Furthermore, alcohol use likely influenced baseline TB status.[38, 58, 56] Under the target trial framework, SEARCH-TB would result in “Emulation failure 2: time zero is set at eligibility but after strategy assignment”.[26] However, we now demonstrate how following the Causal Roadmap facilitates clear definition of the causal estimand with Counterfactual Strata Effects, while avoiding bias due to missing data during estimation and inference.
+To do so, we update our notation to have multiple outcome measures.
 Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”).
-Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes; they are set to NA when not observed.
+Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes.
 The corresponding causal models can be found in Figure 3.
-The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$.
-As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at risk of TB at baseline ($Y_{0}=0$).
-Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline.
+The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$.[39, 40, 1] As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at-risk of TB at baseline ($Y_{0}=0$).
+Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and at endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline.
 For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
-To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$.
-Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at risk at baseline.
+To define the causal estimand in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$.
+Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at-risk at baseline.
 Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline.
-Thereby, this final intervention is a dynamic or personalized one to set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise (e.g., [24, 41, 42, 43]).
+Thereby, this final intervention is a dynamic or personalized.[7, 29, 74, 62] We set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise.
 For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
-Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$.
-Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”[24, 5, 25, 26, 27, 28], which are defined by Nakato et al.
-as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure” [29].
-These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [44, 45, 46, 47].
-For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use.
-Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk at baseline.
-To identify this effect, we re-express the conditional probability as
+Now we can precisely define the causal estimand in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among all at-risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$.
+Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”, which are defined by Nakato et al.
+as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure”.[47] These estimands have previously been specified to evaluate intervention effects in cluster randomized trials.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] To our knowledge, this is their first use in observational studies and with additional missingness on the exposure.
+To facilitate identification, we re-express the conditional probability as
 $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$ (1)
-Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn.
-The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk.
+Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify the denominator and the numerator.
+The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having TB at baseline — under alcohol use $A=a$ and under complete measurement of the exposure and baseline outcome ($\Delta_{A}=\Delta_{Y_{0}}=1)$.
 The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome.
 Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S3.3.
-Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
+Since we are interested the counterfactual probability of being at-risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
+Altogether, this statistical estimand accounts for differential measurement of the exposure and baseline TB status as well as the impact of alcohol use $A=a$ on baseline TB risk.
 In our final causal parameter (Eq.
-1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at endline but not at baseline.
+1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of acquiring TB — under alcohol use $A=a$ and under complete measurement $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$.
 For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values.
 To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following.
-Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
+Among those known to be at-risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
 We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e.
@@
 $\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
-Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at endline is given by
+Putting it all together, our statistical estimand is given by
 $\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
@@
 In the previous section, we introduced a series of causal models and identification results of increasing complexity.
-For the resulting statistical estimands, we could use a singly robust estimation approach, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW) [9, 13].
-Here, we highlight the use of TMLE, which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [11].
-In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation.
-Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [48].
-Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions.
+For the resulting statistical estimands, we could use singly robust estimators, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW).[61, 31] Here, we highlight the use of TMLE, which is a doubly robust procedure and asymptotically efficient under certain conditions.[76] In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation.
+Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners.[75] Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions.
 Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq.
-2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
-TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [11].
-The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [49].
-The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values.
+Appendix S4.1: Simulation Results), we would implement a separate TMLE for the numerator estimand (the joint probability) and the denominator estimand (1- the baseline prevalence) before combining the results.
+TMLE is an asymptotically linear estimator and is normally distributed in the large data limit.[76] The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability.[79] The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values.
 Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve.
 For ratio-type estimands (Eq.
-2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand.
-Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
+Appendix S4.1: Simulation Results), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand.[79, 45] Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals (Appendix S4).
+To evaluate the practical performance of our approach, we conducted a simulation study reflecting the motivating study (Figure 3).
+The simulation design and results are detailed in Appendix S5.
+Briefly, we found that TMLE had negligible bias and achieved nominal confidence interval coverage for the exposure-specific estimands $\Psi(\mathbb{P};a)$, given in Eq.
+Appendix S4.1: Simulation Results for $A=\{0,1\}$, as well as their ratio (i.e.
+the relative effect) and their difference (i.e., the average treatment effect).
+All other approaches exhibited meaningful bias and sub-nominal confidence interval coverage.
 Application to SEARCH-TB
 We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda?
-With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets [21, 22, 23, 50, 38].
-Concretely, the causal model reflected team’s knowledge of study protocol, the study procedures, and the epidemiology of TB in the region.
+With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan.[53, 15, 14, 21, 48, 1] Our causal model reflected the team’s knowledge of the study protocol, study procedures, and epidemiology of TB in the region.
 Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures.
-For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean.
-We conducted influence curve-based inference, accounting for clustering by household (Appendix S4) [51, 26].
+The exposure was self-report of any alcohol use, and the outcome was incident TB infection: conversion from a negative tuberculin skin test to a positive one.[1] The covariates and exposure were measured at the start of the parent SEARCH study (2013-2014), while TB status was measured subsequently in the sub-study (2015-2017).[25, 39, 40] For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean.
+We conducted influence curve-based inference, accounting for clustering by household (Appendix S6).[78, 50]
 In secondary analyses, we considered communities, instead of households, to be the independent unit.
-To examine the sensitivity of our results to alternative estimation approaches, we also implemented IPW for the same statistical estimand, but using parametric regressions to estimate the weights.
-Finally, to examine the impact of handling of our missing data, we took the following “naïve” approach: (1) subset on participants known to be at risk of TB at baseline ($Y_{0}=0$); (2) further subset on participants with measured exposures ($\Delta_{A}=1$) and measured outcomes at endline ($\Delta_{Y_{1}}=1$), and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)\mid Y_{0}=0]$.
+To examine the sensitivity of our results to alternative estimation approaches, we implemented two singly robust approaches for the final statistical estimand (Eq.
+Appendix S4.1: Simulation Results): G-Computation using parametric regressions to estimate the iterated conditional expectations and IPW using parametric regressions to estimate the weights.[8] Finally, to examine the impact of our handling of missing data, we took the following “naïve” approach: (1) subset on participants with complete data $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$; (2) further subset on participants known to be at-risk at baseline $Y_{0}=0$, and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)]$ among the remaining subset of participants.
 This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3).
-Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding [52, 53, 54].
-For statistical inference, both IPW and the naïve approach accounted for clustering by household.
-In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [38].
-As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S4.
-In contrast, IPW relying on parametric assumptions resulted in a smaller association and confidence intervals overlapping the null: RR=1.13 (95%CI: 1.00-1.27).
-Finally, after restricting to participants who tested negative at baseline, responded to questions about alcohol use, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
-Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at risk at baseline and with measured exposures and outcomes at endline.
+Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding.[64, 41, 17] For statistical inference, G-computation, IPW, and the naïve approach accounted for clustering by household.
+In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59).[1] As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S6.
+In contrast, G-computation resulted in a larger association with large confidence intervals, while IPW resulted in a smaller association and confidence intervals overlapping the null: RR=1.58 (95%CI: 1.36-1.83) and RR=1.13 (95%CI: 1.00-1.27), respectively.
+Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
+Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “G-Comp” with G-computation, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at-risk at baseline and with measured exposures and outcomes at endline.
 Discussion
-We presented causal models, causal parameters, and identification results for a series of prospective studies with increasing levels of missingness.
+We presented causal models, causal estimands, and identification results for a series of prospective studies with increasing levels of missingness.
+We reviewed how the Causal Roadmap could be applied to address missing outcomes and extended to address missing exposures.
+We then motivated the use of Counterfactual Strata Effects to precisely define causal estimands when the focus population is subject to missingness and/or impacted by the exposure.
+Specifically, we considered the common scenario where there is missingness on the outcome at baseline — determining who is at-risk of acquiring the outcome by endline.
+We further allowed for the exposure to influence baseline outcome risk.
+We demonstrated that while these real-world complications would result in “failures” in the target trial emulation framework,[26] our Roadmap approach facilitates definition and identification of the corresponding causal estimand.
 For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands.
-Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work.
-Using TMLE to flexibly account for confounding and missingness, we found a 49% relative increase in the risk of incident TB infection associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59).
-For the same statistical estimand but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27).
-Finally, the naïve approach to condition on the participants with baseline TB risk and measured exposures/outcomes yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
-There are several advantages to our framework for handling missing exposures and outcomes when evaluating the causal effects.
-Unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data [4], our approach focuses our efforts on the portions of the observed data distribution relevant for the statistical estimand.
-Furthermore, our approach uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (sometimes called an “available-case analysis”) [4].
+We emphasize that Counterfactual Strata Effects are distinct from Principal Strata Effects.[16, 18, 19, 51] For simplicity, suppose that we did not have missingness, and define $Y_{0}^{a}$ and $Y_{1}^{a}$ as the counterfactual outcome at baseline and endline under exposure $A=a$.
+Then in SEARCH-TB, principal stratification could be applied to define the effect of alcohol on incident TB among the unobservable subset of participants who would have always been at-risk of TB at baseline regardless of their alcohol use: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=Y_{0}^{0}=0)$ versus $\mathbb{P}(Y_{1}^{0}=0\mid Y_{0}^{1}=Y_{0}^{0}=0)$.
+Instead, Counterfactual Strata Effects enable us to define the effect of alcohol on incident TB among the counterfactual population of persons at-risk at baseline under alcohol use $A=a$: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=0)$ versus $\mathbb{P}(Y_{1}^{0}=1\mid Y_{0}^{0}=0)$.
+The at-risk population is exposure-specific, and our estimand captures the effect of alcohol use on prevalent TB at baseline and on incident TB by endline.
+Of course, we could also directly examine the effect on prevalent TB at baseline: $\mathbb{P}(Y_{0}^{1}=1)$ versus $\mathbb{P}(Y_{0}^{0}=1)$.
+In addition to incidence studies,[40, 50] Counterfactual Strata Effects commonly arise when studying care cascades: the series of steps from screening, through diagnosis and treatment, to control.[7, 54, 2, 25, 5, 4, 47] Indeed, these cascades are a series of conditional probabilities — each subject to missingness and each potentially impacted by the exposure or intervention strategy.
+There are several advantages to our framework for handling multi-source missingness.
+As previously discussed, our approach avoids a common “failure” arising when aiming to emulate at target trial.[26] Furthermore, unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data,[46] our approach focuses our efforts on the portion of the observed data distribution relevant for the statistical estimand (Appendix S7).
+Our approach also uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (Figure 4; Appendix S.5).
 Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
-A major limitation to this work is that we did not address missingness on the confounders.
-Some approaches include mean/median imputation or including missingness indicators in the adjustment set.
-For further discussion on missing confounders, we refer to the comprehensive work of Williamson et al.
-[55].
-We also focused on cross-sectional or prospective studies.
+There are several limitations to our work.
+First, we focused on cross-sectional or prospective studies.
 Thus, we did not cover scenarios where the outcome impacts the measurement of other variables.
-Such scenarios would arise in case-control studies and have been addressed in prior literature (e.g., [56, 57, 58, 59, 54]).
-There are other limitations to our work.
-First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health.
-Second, in our real-data demonstration, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [60, 61, 62].
-Future work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants.
-Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness.
-In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses [63, 64].
-Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [21, 65, 23, 50].
-Overall, our goal was to demonstrate how an established framework for estimating causal effects with censored outcomes could be extended to settings with missingness on the exposure, the outcome at baseline, and the outcome at endline.
-To do so, we provided a series of causal models with increasing complexity and discussed the identification assumptions for each.
-For estimation and inference, we focused on TMLE, a doubly-robust approach using machine learning and data from all participants (including those with missing exposures or outcomes).
-Our work was motivated by the investigation of the effect of alcohol use on incident TB infection [38], and our re-analysis demonstrated the real-world consequences of our approach.
-Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
-Conflicts of Interest: The authors report no conflict of interest in this work.
-Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved.
-We also thank Dr.
-Diane Havlir and Dr.
-Maya L.
-Petersen, who together with Dr.
-Moses R.
-Kamya are the MPIs of the SEARCH collaboration.
-Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done.
-Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
+Such scenarios would arise in case-control studies and have been addressed in prior literature.[65, 63, 86, 34, 17, 57] Second, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health.
+Third, we did not consider multiple imputation, which is a common approach for missing data.[66] However, foundational work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants.
+Indeed, Tompsett et al.
+suggested that multiple imputation cannot be used in settings with missingness on the focus population (termed “missing eligibility data”) and missingness on the exposure.[71] Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness.
+In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses.[36, 13] Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question, and appropriately addresses many real-world complications.[53, 15, 14, 21, 48]
 Supplementary Materials for “Causal Inference with Missing Exposures and Missing Outcomes”
 Contents:
-  • Appendix S1: More on the classic point-treatment problem
+  • Appendix S1: The classic point-treatment problem
+  • Appendix S2: Missing confounders
   • Appendix S3: Proofs
+  • Appendix S4: More on the Delta method
+  • Appendix S5: Finite sample simulation study
   • Appendix S6: Accounting for outcome dependence
-Appendix S1: More on the classic point-treatment problem
+  • Appendix S7: Missingness DAGs
+Appendix S1: The classic point-treatment problem
 We consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close.
@@
 Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$.
-Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$.
+Then our causal estimand $\mathbb{E}(Y^{a})$ is the expected counterfactual outcome if all had exposure-level $A=a$.
 In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
@@
 To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$.
-This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$ [9, 10, 11, 12].
-Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$.
+This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$.[61, 52, 76, 28] Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$.
 In Figure 5B and C, we show two causal models where this assumption would hold.
@@
 We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model.
-Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [9].
-Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
-In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation [21, 22, 23, 66, 65] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
+Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$.[61] Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
+In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation.[53, 15, 21, 82, 14] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
+Appendix S2: Missing Confounders
+Throughout the main text, we also assume that the covariates are completely measured.
+While missingness on covariates is common, a detailed discussion is beyond the scope of this manuscript, which focuses on defining, identifying, and estimating causal effects in settings with missing exposures and missing outcomes.
+Here, we provide a brief review of methods to handle missing confounders.
+Common approaches include dropping covariates subject to missingness, conducting a complete (available) case analysis, including a missingness indicator, and using multiple imputation.[20, 83, 81, 57] Excluding covariates subject to missingness will induce bias if those covariates are confounders.
+A complete case analysis, where observations with missing data are excluded, is guaranteed to reduce power and will be biased unless missingness is completely at random.
+In the missing indicator method, a dummy variable is included to indicate the covariate is measured/missing.
+Recently, the missingness indicator method was shown to perform similarly to multiple imputation (MI) in many settings.[83] MI uses observed variables to predict the values of missing ones and has traditionally relied on strong modeling assumptions, although there is a growing interest in using machine learning.[67, 35, 10]
+Williamson et al.
+provide an excellent discussion and evaluation of more modern methods for missing confounders, including doubly robust approaches of generalized raking and targeted minimum loss-based estimation (TMLE).[81] These approaches are more robust to model misspecification and are efficient in certain scenarios.
+Williamson et al.
+also provide guidance on selecting the approach to handle missing confounders based on the bias-variance trade-off.[81] Recently, Qiu et al.
+extended this work and showed TMLE, a plug-in estimator, had superior performance to generalized ranking in finite samples.[57]
 Appendix S3: Proofs
@@
 $\text{by }Y^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y^{*}\perp A\mid\Delta_{A}=1,L_{0}$
-$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
+$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
 $\text{by }Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$
-$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},\Delta_{A}=1,A=a,L_{0}=l_{0})\times$
-$\quad\mathbb{P}(L_{1}=l_{1}\mid\Delta_{a}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
+$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
+$\quad\mathbb{P}(L_{1}=l_{1}\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
 $=$ $\mathbb{E}\bigg\{\mathbb{E}\big[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})\big]\bigg\}$
@@
 Appendix S2.3: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 3 in the main text)
-Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as
+Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
+Recall that we defined the target parameter for this section using Counterfactual Strata Effects:[3, 54, 5, 4, 50, 55, 22, 25, 40]
 $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$
@@
 in addition to the positivity assumptions for the denominator.
+Appendix S4: The Delta Method for Ratio-Estimands
+The Delta method is commonly applied to derive influence curve-based inference for asymptotically linear estimators.[79] Recall our exposure-specific estimand is given by the following ratio:
+$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (3)
+Following the notation of Moore and van der Laan,[45] we denote the numerator with $\mu_{1}$ and the denominator with $\mu_{0}$.
+Let $IC_{1}$ and $IC_{0}$ be the corresponding influence curves.
+Then based on the Delta method, the influence curve for $log[\Psi(\mathbb{P};a)]$ is given by $IC(a)=1/\mu_{1}\times IC_{1}-1/\mu_{0}\times IC_{0}$.
+We obtain a variance estimate with the sample variance of $IC(a)$ scaled by sample size $N$.
+Using that variance estimate, we conduct hypothesis testing and create 95% confidence intervals before exponentiating to transform back to the original scale.
+Now consider the relative association:
+$RR=\frac{\Psi(\mathbb{P};1)}{\Psi(\mathbb{P};0)}=\frac{\mu_{1}/\mu_{0}}{\mu_{3}/\mu_{2}}$ (4)
+with corresponding subscripts for the influence curves of the TMLEs. Then the influence curve for the $log(RR)$ is given by
+$IC_{RR}=(-1/\mu_{0})\times IC_{0}+(1/\mu_{1})\times IC_{1}+(1/\mu_{2})\times IC_{2}-(1/\mu_{3})\times IC_{3}$ (5)
+We obtain inference as above.
+Appendix S5: Simulation Study
+We conducted a simulation study designed to mimic the causal model given in Figure 3 in the main text. The data generating process is as follows where $\text{expit}(x)=(1+e^{-x})^{-1}$.
+  • Baseline covariates: $L_{0,1}\sim\text{Bernoulli}(0.5)$, $L_{0,2}\sim\text{Uniform}(0,1)$, and $L_{0,3}\sim\text{Uniform}(0,1)$.
+  • Measurement indicator for the exposure: $\Delta_{A}\sim\text{Bernoulli}(\text{expit}(1.7+0.4L_{0,1}-0.5L_{0,3}))$
+  • Exposure: $A\sim\text{Bernoulli}(\text{expit}(-0.3+1.5L_{0,1}-1.2L_{0,2}))$, observed only when $\Delta_{A}=1$.
+  • Measurement indicator for the baseline outcome: $\Delta_{Y_{0}}\sim\text{Bernoulli}(\text{expit}(2.0+0.3L_{0,1}-0.4L_{0,2}-0.3A+0.2\Delta_{A}))$
+  • Baseline outcome: $Y_{0}\sim\text{Bernoulli}(\text{expit}(-2.6+1.2L_{0,1}+0.8L_{0,2}-0.5L_{0,3}+0.7A+0.3A\cdot L_{0,1}))$, observed only when $\Delta_{Y_{0}}=1$.
+  • Time-varying covariates: $L_{1,1}=\text{clip}(\mathcal{N}(0.3A+0.2L_{0,1}+0.15L_{0,2},\ 0.25^{2}),0,1)$ and $L_{1,2}\sim\text{Bernoulli}(\text{expit}(-0.3+1.2A+0.8L_{0,1}-0.6L_{0,3}+0.5Y_{0}))$
+  • Measurement indicator for the endline outcome: $\Delta_{Y_{1}}\sim\text{Bernoulli}(\text{expit}(1.1+5.5L_{1,1}^{2}A-5.0L_{0,3}L_{1,1}L_{1,2}-4.0Y_{0}A+0.5L_{0,1}))$, among those with $\Delta_{Y_{0}}=1$ and equal to 0 otherwise.
+  • Endline outcome: $Y_{1}\sim\text{Bernoulli}(\text{expit}(-2.0+5.0A\cdot L_{1,1}^{2}-4.5L_{0,3}^{2}L_{1,2}+1.5L_{0,1}+1.2A))$, observed only when $\Delta_{Y_{1}}=1$.
+We calculated the true value of the exposure-specific estimands $\Psi(\mathbb{P};1)$ and $\Psi(\mathbb{P};0)$ as well as the relative and absolute effects by generating $10^{6}$ counterfactuals and taking the relevant means, ratios, and differences, respectively.
+We implemented TMLE, G-computation, and inverse probability weighting (IPW) for the exposure-specific estimands and the effects.
+Mirroring the real-data analysis, we used Super Learner within TMLE to combine predictions from generalized linear models, the mean, and multivariate adaptive regression splines with and without covariate screening.
+We used parametric regressions in the singly robust approaches.
+Also mirroring the real-data analysis, we implemented a “Naïve-adjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then implemented TMLE with adjustment for confounding.
+For completeness in the simulations, we also implemented a “Naïve-unadjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then take the empirical mean outcome within exposure groups.
+Influence curve-based inference was used for variance estimation and confidence interval construction.
+We ran 1000 simulation studies at a sample size of $N=20000$ to reflect the motivating example.
+Appendix S4.1: Simulation Results
+The true values of the estimands were as follows: the expected outcome under the exposure $\Psi(\mathbb{P};1)=0.510$, the expected outcome under no exposure $\Psi(\mathbb{P};0)=0.177$, the risk ratio $RR=2.879$, and the risk difference $RD=0.333$.
+Table 1 provides results in terms of bias (average deviation between the point estimate and truth), empirical standard deviation (standard deviation [SD] of the point estimate across simulation replicates), the ratio of bias to empirical SD, and 95% confidence interval coverage (proportion of calculated intervals containing the truth).
+We report the bias-to-SD ratio as a diagnostic for the asymptotic behavior of our estimators: since the standard deviation shrinks at rate $\sqrt{N}$ while systematic bias remains fixed, a non-negligible bias-to-SD ratio indicates performance will deteriorate as sample size grows.
+For all estimands, TMLE with Super Learner achieved nominal coverage, had negligible bias, and had bias-to-SD ratios well below 0.2.
+In contrast, G-computation and IPW only achieved nominal coverage for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, but not the other estimands.
+Indeed, the coverage of G-computation was 93.2% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 91.6% for the risk ratio, and 92.1% for the risk difference.
+The coverage of IPW was only 12.3% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 19.7% for the risk ratio, and 58.9% for the risk difference.
+Table 1: Over 1000 simulation iterations, estimator performance for TMLE, G-computation (“G-comp”), inverse probability weighting (IPW), and the naïve approach with and without confounding adjustment.
+Estimand Estimator Bias Emp SD Bias/SD Coverage
+$\Psi(\mathbb{P};1)$ TMLE 0.0014 0.0091 0.1560 0.9440
+$\Psi(\mathbb{P};1)$ G-comp 0.0053 0.0083 0.6435 0.9500
+$\Psi(\mathbb{P};1)$ IPW -0.0029 0.0089 -0.3231 0.9690
+$\Psi(\mathbb{P};1)$ Naïve-adjusted 0.0565 0.0088 6.4347 0.0000
+$\Psi(\mathbb{P};1)$ Naïve-unadjusted 0.1089 0.0080 13.5559 0.0000
+$\Psi(\mathbb{P};0)$ TMLE 0.0003 0.0058 0.0578 0.9530
+$\Psi(\mathbb{P};0)$ G-comp -0.0024 0.0054 -0.4546 0.9320
+$\Psi(\mathbb{P};0)$ IPW 0.0179 0.0063 2.8376 0.1230
+$\Psi(\mathbb{P};0)$ Naïve-adjusted 0.0102 0.0060 1.7023 0.6290
+$\Psi(\mathbb{P};0)$ Naïve-unadjusted -0.0151 0.0052 -2.9191 0.0380
+Risk Ratio TMLE 0.0056 0.1068 0.0525 0.9510
+Risk Ratio G-comp 0.0733 0.1012 0.7247 0.9160
+Risk Ratio IPW -0.2765 0.0953 -2.9025 0.1970
+Risk Ratio Naïve-adjusted 0.1484 0.1054 1.4071 0.7240
+Risk Ratio Naïve-unadjusted 0.9454 0.1316 7.1833 0.0000
+Risk Difference TMLE 0.0011 0.0107 0.1012 0.9460
+Risk Difference G-comp 0.0078 0.0097 0.7978 0.9210
+Risk Difference IPW -0.0208 0.0109 -1.9127 0.5890
+Risk Difference Naïve-adjusted 0.0463 0.0104 4.4549 0.0100
+Risk Difference Naïve-unadjusted 0.1240 0.0095 13.0900 0.0000
+We emphasize that TMLE, G-computation, and IPW rely on the same identification result.
+Therefore, differences in their performance are due to differences in statistical estimation.
+To explore further, Table 2 provides their performance for the numerator and denominator of our ratio-estimand:
+$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$
+TMLE had negligible bias for all components; its bias-to-SD ratio remained below 0.2.
+In contrast, G-computation exhibited meaningful bias, as indicated by the bias-to-SD ratio, for all components except for the numerator under $A=1$.
+The numerator/denominator decomposition identifies the source IPW’s poor performance: the numerator for $A=0$ is substantially biased with a bias-to-SD ratio of $2.79$.
+This bias propagates directly into meaningful bias and poor coverage for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, the risk ratio, and the risk difference.
+Table 2: Further details on estimator performance with TMLE, G-comp, and IPW for the exposure-specific components of the ratio-estimands.
+Estimator Exposure Component Truth Mean Est Bias Emp SD Bias/SD
+TMLE $A=1$ Numerator 0.3593 0.3602 0.0009 0.0071 0.1274
+TMLE $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0346
+TMLE $A=0$ Numerator 0.1499 0.1501 0.0002 0.0050 0.0419
+TMLE $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0044 -0.0957
+G-comp $A=1$ Numerator 0.3593 0.3604 0.0011 0.0066 0.1686
+G-comp $A=1$ Denominator 0.7047 0.6996 -0.0051 0.0054 -0.9529
+G-comp $A=0$ Numerator 0.1499 0.1473 -0.0026 0.0046 -0.5549
+G-comp $A=0$ Denominator 0.8464 0.8435 -0.0029 0.0044 -0.6663
+IPW $A=1$ Numerator 0.3593 0.3571 -0.0021 0.0069 -0.3099
+IPW $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0344
+IPW $A=0$ Numerator 0.1499 0.1649 0.0151 0.0054 2.7942
+IPW $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0045 -0.0950
+Now we consider the naïve approaches, which are targeting a different (incorrect) estimand.
+Therefore, there was a substantial “causal gap”, even when adjusting for confounding.[53, 15, 14, 21, 48] By simply subsetting on participants with complete data and baseline risk, the naïve approaches failed to appropriately account for differential missingness.
+As shown in Table 1, the naïve approaches exhibited meaningful bias for all estimands.
+The coverage of naïve approach with confounding adjustment was 0% for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, 62.9% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 72.4% for the risk ratio, and 1% for the risk difference.
+Finally, the coverage of naïve approach without confounding adjustment was <<5% for all estimands.
+Altogether, our simulation results support the robustness of TMLE with Super Learner in settings resembling our applied analysis.
+Our simulation results further suggest the practical differences between TMLE, G-computation, and IPW in SEARCH-TB are due to TMLE’s flexible adjustment for confounding and differential missingness.
+Again, all three approaches are targeting the same estimand.
+Therefore, the simulation and applied results highlight how theoretical properties (e.g., double robustness) translate into practical impact.
 Appendix S6: Accounting for outcome dependence
@@
 This dependence should be reflected in the corresponding causal model (e.g., [23, 24, 32, 6]).
-By following the Causal Roadmap or a similar framework for causal inference [21, 71], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [51, 70, 25, 26, 72]).
+By following the Causal Roadmap or a similar framework for causal inference,[53, 27] we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [78, 6, 4, 50, 49]).
 Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households.
@@
 The exact form of the causal model and identification result will depend on the application.
-Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [51, 26, 72].
+Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit.[78, 50, 49]
 If clustering is present, estimation and inference must be adjusted.
@@
 Concretely, participants in a given cluster are all assigned to the same sample-split.
-Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [18].
-Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.
-Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$.
+Second, variance estimation must account for clustering.
+Following Schnitzer et al.,[68] let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$.
+Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.[68] Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$.
 We then proceed with variance estimation using the cluster-level influence curve.
 This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.
+(As detailed [9], a slightly different formulation is needed if the causal estimand weights clusters, versus individuals, equally.)
+Appendix S7: Missingness DAGs
+There is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”).[43, 44] In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link.[46] They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures.[46, 85] Holovchak et al.
+extended this work for the effects of longitudinal exposures, while questioning the practical utility of the m-DAG approach.[30] Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis”.[30] As an alternative, we demonstrate how the Causal Roadmap, an established framework for causal inference, can used to define, identify, and estimate causal estimands, including the average treatment effect under missing exposures and Counterfactual Strata Effects.